# Technical Note: Vendor-Specific Approach for Standardized Uptake Value Calculation

### v1

@@ 5 Conclusions
The proposed vendor-specific SUV calculation strategy improves PET imaging accuracy by addressing key discrepancies in handling PET-related DICOM attributes across different vendors. Our method effectively mitigates errors caused by the misalignment of series time and acquisition time, issues commonly found in vendor-neutral approaches. Additionally, we have implemented this vendor-specific strategy in the open-source software Z-Rad, providing a ready-to-use tool for accurate SUV calculation from DICOM files of different PET scanner vendors.
@@ 6 Figures and Tables
Data Source Vendor Manufacturer Model Name (0008,1090) Units (0054,1001) No. of PET Images
autoPET II 10 Siemens Biograph mCT, mCT Flow, Biograph 64 BQML (100%) 1014
USZ GE Discovery MI, Discovery 690, Discovery STE, Discovery 710, Discovery RX BQML (100%) 522
USZ Philips Allegro Body, GEMINI TF 64, GEMINI TF Big Bore, or not specified BQML(57%), CNTS (43%) 28
@@
Table 2: DICOM attributes
@@
Decay Correction DateTime (0071, 1022) Siemens -

### v2

@@ 4 Conclusions
Our proposed vendor-specific SUV calculation strategy substantially improves PET quantitative analyses by accommodating vendor-specific DICOM attributes and correcting for discrepancies in Series Time (0008,0031) and Acquisition Time (0008,0032). These adjustments resolve the common pitfalls associated with the QIBA approaches, particularly in cases where Decay Correction Date Time attributes are either missing or misaligned.
5 Acknowledgments
This study was supported by Comprehensive Cancer Center Zurich (C3Z Precision Oncology Funding Program, OMD project). We would also like to thank GE and Siemens Healthineers for their helpful input regarding SUV definitions specific to their respective systems.
6 Conflict of Interest
The department of Radiation Oncology, University Hospital Zurich has teaching and research agreements with Siemens Healthineers. Stephanie Tanadini-Lang received travel funds from Siemens Healthineers. The remaining authors declare no conflict of interest.
@@ 7 Figures and Tables
Data Source Vendor Manufacturer Model Name (0008,1090) Units (0054,1001) No. of PET Series
autoPET II autopet_dataset, gatidis_kuestner_2022 Siemens Healthineers Biograph mCT, mCT Flow, Biograph 64 BQML (100%) 1014
External
and internal data
GE Discovery MI, Discovery 690, Discovery STE, Discovery 710, Discovery RX BQML (100%) 522
External data,
ACRIN 6685Kinahan2019
Philips Allegro Body, Guardian Body, GEMINI TF 16, GEMINI TF 64, GEMINI TF Big Bore, or not specified BQML(70%), CNTS (30%) 71
IBSI II vallieres_2015_tcia, vallieres_2015_pmb, clark_2013_tcia GE Discovery ST GML (100%) 51
@@
Table 2: *
DICOM attributes used for SUV calculation.
@@
Decay Factor (0054,1321) No The decay factor that was used to scale this frame.
Frame Reference Time (0054,1300) No The time that the pixel values in the image occurred. Frame Reference Time is the offset, in msec, from the Series reference time.
Actual Frame Duration (0018,1242) No Elapsed time for data acquisition in msec.
@@
Decay Correction DateTime (0071,1022) Siemens Healthineers -